OpenStudy (anonymous):

Can someone please help A cylinder is inscribed in a right circular cone of height 2.5 and radius (at the base) equal to 5.5. What are the dimensions of such a cylinder which has maximum volume?

OpenStudy (anonymous):

V=pi*r^2*h v=3.14*5.5^2*2.5 v=3.14*30.25*2.5 v=237.4625

OpenStudy (anonymous):

just use formula.

OpenStudy (anonymous):

hope it helps you.

OpenStudy (anonymous):

How do i find Height and Radius from this ?

OpenStudy (anonymous):

It gives you the radius and the height in the question. Reread it. You have to find the volume which is V. Don't let the question confuse you. It give you the height which is 2.5 then it gives you the radius (at the bass) 5.5. Now all you have to do is fill in the blanks of this formula V=pi*r^2*h which is the answer i gave you above. The question is trying to confuse you just pay close attention. message me if you need more help. If you want!

OpenStudy (anonymous):

Thank You! :)

OpenStudy (anonymous):

No problem.